the product of two consecutive positive intergers is 41 more than their sums find the intergers
Answer by checkley77(12844)

x(x+1)=x+x+1+41
x^2+x=2x+42
x^2+x-2x-42=0
x^1-x-42=0
(x-7)(x+6)=0
x-7=0
x=7 ans
7+1=8 ans.
Proof:
7*8=*7+8+41
56=15+41
56=56